OpenStudy (anonymous):

Which value for x makes the open sentence true? 8 + 3 • x = 22 + x2 A. 5 B. 4 C. 3 D. 2

OpenStudy (anonymous):

billguy

OpenStudy (anonymous):

I

OpenStudy (anonymous):

I WILL MEDAL ANYONE

OpenStudy (anonymous):

You just have to substitute the given value in for x and perform the operation. If the numbers on each side of the equal sign equal each other after the substitution then the statement is true

OpenStudy (anonymous):

I need Help solving it

OpenStudy (anonymous):

For example A gives you 8 + 3 (5) = 22 + 2(5) which is 23=32, this is not true so it isn't A

OpenStudy (anonymous):

Hmm, none of these seem to add up. Did you copy the problem correctly? Is the last term supposed to be X to the power of 2 or X times 2?

OpenStudy (anonymous):

8 + 3 • x = 22 + x2

OpenStudy (anonymous):

Choose the value for x that makes the open sentence true. 8 + 3 • x = 22 + x2 (Points : 1) 5 4 3 2

OpenStudy (anonymous):

There you are

OpenStudy (anonymous):

Choose the value of x that makes the open sentence true. 40 – x2 < 2x + 15 (Points : 1) 2 3 4 5

OpenStudy (anonymous):

Choose the value of x that makes the open sentence true. 40 – x2 < 2x + 15